The front 1.20 m of a 1,250-kg car is designed as a "crumple zone" that collapses to absorb the shock of a collision.


(a) If a car traveling 26.0 m/s stops uniformly in 1.20 m, how long does the collision last?

___ s


(b) What is the magnitude of the average force on the car?

 ___N


(c) What is the magnitude of the acceleration of the car? Express the acceleration as a multiple of the acceleration of gravity.

 ___ g

1
Expert's answer
2021-10-02T14:56:01-0400

Solution.

"L=1.20m;"

"m=1250kg;"

"v_0=26.0m\/s;"

"v=0m\/s;"

a) "L=\dfrac{v_0+v}{2}t\implies t=\dfrac{2L}{v+v_0};"

"t=\dfrac{2\sdot1.20}{26.0}=0.92s;"

b) "F=ma;"

"a=\dfrac{v-v_0}{t}" ;

"a=\dfrac{0-26.0}{0.92}=-28.26 m/s^2;"

"F=1250\sdot28.26=35326N;"

c)"g=9.8m\/s^2;\na=2.88g;"

Answer: "a) t=0.92s;"

b)"F=35326N;"

c)"a=2.88g."